OpenStudy (anonymous):

It is easy to check that for any value of x=y*e^(C*y+1), the function is solution of equation y'=y/(x(ln(x)-ln(y)). Find the value of for which the solution satisfies the initial condition y(1)=1 . c=?

OpenStudy (anonymous):

at x=1 y=1 \[x=y*e ^{Cy+1}\] \[1=1*e ^{C.1+1}\] \[1=e ^{C+1}\] \[ln(1)=ln(e ^{C+1})\] \[0=C+1\] \[C=-1\]